OpenStudy (anonymous):

The distance of a train from a station, varies directly with the time, t. If d = 200 miles when t = 4 hours, find d when t = 6.

OpenStudy (mayankdevnani):

Any ideas? @balokkk

OpenStudy (anonymous):

not at all

OpenStudy (mayankdevnani):

ohkk

OpenStudy (mayankdevnani):

\[\huge \bf d=k \times t\] where d=distance , t=time and k=constant value. If d=200 miles and t=4 hrs, \[\huge \bf200=k \times 4\] so, \[\huge \bf k=\frac{200}{4}=50\]

OpenStudy (mayankdevnani):

If time=6 hrs and d=? \[\huge \bf k=\frac{d}{t}\] \[\huge \bf 50=\frac{d}{6}\] solve for d .

OpenStudy (mayankdevnani):

can you solve it? @balokkk

OpenStudy (anonymous):

d=300?

OpenStudy (mayankdevnani):

yeah. you are correct. @balokkk

OpenStudy (anonymous):

thanks you

OpenStudy (mayankdevnani):

welcome :)